Circular No. 7895

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J0929–314 A. J. Castro-Tirado, Instituto de Astrofísica de Andalucía, Consejo Su-

perior de Investigaciones Científicas (IAA-CSIC), Granada; A. Caccianiga, Osservatorio Astronomico di Brera, Milano; J. Gorosabel, IAA-CSIC; P. Kilmartin, University of Canterbury; P. Tristram and P. Yock, University of Auckland; C. Sánchez-Fernández, Laboratorio de Astrofísica Espacial y Física Fundamental, Instituto Nacional de Técnica Aeroespacial, Madrid; and M. E. Alcoholado-Feltström, Sociedad Malagueña de Astronomía, Málaga, communicate: “We have observed the optical counterpart of the x-ray transient XTE J0929–314 (IAUC 7888, 7889, 7893) with the European Southern Observatory 3.6-m telescope at La Silla. Two 1200-s spectra (350–800 nm) were obtained on May 6.96 and 7.96 UT. The combined spectrum shows emission lines from the C III-N III (464.0–465.0 nm) blend (EW = 0.14 nm for N III) and Hα (656.3 nm, EW = 0.14 nm). These lines are superposed on a blue continuum and are typical of soft-x-ray transients in outburst. Monitoring with the 0.6-m telescope (+ MOA camera) at Mt. John Observatory reveals that the counterpart has not changed in brightness by more than 0.1 mag since May 1.6. Additional spectroscopy and photometry during the outburst will be highly valuable.”

V2116 OPHIUCHI F. Jablonski, D. Cieslinski, and J. Braga, Instituto Nacional de Pes-

quisas Espaciais, São José dos Campos; and M. G. Pereira, Universidade Estadual Feira de Santana, report that they have detected optical pulsations from V2116 Oph, the visible/infrared counterpart to GX 1+4, in CCD Cousins-R observations obtained on Mar. 29.21–29.35 (1.6-m telescope) and on Apr. 4.28–4.34 and 5.20–5.34 UT (0.6-m telescope) at Laboratório Nacional de Astrofísica in southern Brazil (cf. Jablonski et al. 1996, Ap.J. 482, L171). The heliocentric period of the pulsations was 137.67 ± 0.07 s on Mar. 29 and 137.3 ± 0.3 s on Apr. 4 and 5, with pulsed semiamplitudes of 2.5, 2.0, and 0.6 percent, respectively. These are the longest values for the rotational period in this pulsar since its discovery in 1970, and the period derivative calculated with respect to the BATSE measurement on 2000 May 26

(www.batse.msfc.nasa.gov/batse/pulsar/data/sources/gx1p4.html)

is +1.63 s/yr (significantly less than the trend of +2 s/yr observed since the beginning of the BATSE era). The optical brightness of V2116 Oph was R = 15.83–16.22, close to the median value for this object.

2002 May 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green